OpenStudy (anonymous):

hello is there anyone to help me solve qustions related to complex number class 11 reply

OpenStudy (anonymous):

What is the question Kapuria?

OpenStudy (anonymous):

the qustion is simlify i35+1/i35

OpenStudy (anonymous):

is it : \(\large\cfrac{i^{35} + 1}{i^{35}} \) ?

OpenStudy (anonymous):

heh noo itss written seprately i35 + 1/i35

OpenStudy (anonymous):

Ok. \(\large i^{35} + \cfrac{1}{i^{35}} \) . Can you find the value for \(\large i^{35} \) ?

OpenStudy (anonymous):

Divide 35 by 4. You get remainder as 3 . So \(\large i^{35} = i^3 \) . As \(\large i^3 = i * i * i = -1 *i = -i \) So : \(\large i^{35} = -i\)

OpenStudy (anonymous):

Plug-in the value of \(\large i^{35} \) and solve it further.

OpenStudy (campbell_st):

well here is a simple thing... use a common denominator \[\frac{i^{35} \times i^{35}}{i^{35} }+ \frac{1}{i^{35}} \] so you have \[\frac{i^{70} + 1}{i^{35}}\] just need to evaluate \[i^{70}\] its not to difficult

OpenStudy (campbell_st):

so when you simplify it you get \[\frac{(i^2)^{35} + 1}{i^{35}}\] substitute \[i^2 = -1\] and work from there

OpenStudy (anonymous):

Will it be better to let it be in the form of \(\large i^{35} + \cfrac{1}{i^{35}} \) \(i^{35} = -i \) It will be easy to move on and simplify.

OpenStudy (anonymous):

tell me the final answer i hve then i34 i + 1/i34 i then wht is the nxt step after thiss

OpenStudy (anonymous):

@kapuria Please look the above posts and try again.

OpenStudy (anonymous):

kkkkkkk

OpenStudy (anonymous):

ans is coming -i is this corrct

OpenStudy (anonymous):

Hmm, no! See, \(\large i^{35} = -i \) , right?

OpenStudy (anonymous):

yess then

OpenStudy (anonymous):

I have : \(\large i^{35} + \cfrac{1}{i^{35}} = -i + \cfrac{1}{-i} = -i + (-i) = - i - i = -2i\)

OpenStudy (anonymous):

kkkkkkk got it thnksss

OpenStudy (anonymous):

You're welcome. :)

OpenStudy (anonymous):

if u donot mind may i ask u 4 more qustions also

OpenStudy (anonymous):

Oh k wait @kapuria

OpenStudy (anonymous):

First of all , I apologize for my mistake. The answer given by me is wrong.

OpenStudy (anonymous):

See, I had : \(-i + \cfrac{1}{-i} \) , \(-i + \cfrac{1}{-i} \times \cfrac{-i}{-i} \) [What I did was just, I multiplied \(\cfrac{1}{-i} \) by \(\cfrac{-i}{-i}\) ]

OpenStudy (anonymous):

Now, \(-i + \cfrac{-i}{-i * -i } \) = \(-i + \cfrac{ -i}{(i^2)} \) \(-i + \cfrac{-i}{-1} \) = \(-i + i = 0\)

OpenStudy (anonymous):

So the answer is 0. [Sorry for my mistake again]

OpenStudy (anonymous):

kkkkkkk r u sure now

OpenStudy (anonymous):

Yep . Very sure. :)

OpenStudy (anonymous):

the qustion is find value of x and y (1+i) y ka squre + (6+i) = (2+i)x solve plzz

OpenStudy (anonymous):

Please post this as a new question. Some one will surely help you.